# OrbitZoo: Multi-Agent Reinforcement Learning Environment for Orbital Dynamics

### v1

@@ 3.4 Multi-Agent Coordination and Scalability
In MARL for orbital dynamics, agents often face partially observable environments with decentralized policies. Challenges include coordinating tasks among agents, avoiding interference, and scaling solutions to larger constellations. Recent works address these challenges by using C++ engines [54], or CPU/GPU parallelization for faster agent training [40] and step/propagation time [30]. The EPyMARL library [37] builds on PyMARL to enable flexible multi-agent configurations, allowing satellite collaboration. However, MARLlib [24], which integrates with PettingZoo environments like OrbitZoo, may be more suitable, as it offers more MARL algorithms and neural network architectures (e.g., recurrent and convolutional).
Centralized critics and techniques like generalized advantage estimation (GAE) [41] improve training stability and efficiency for such environments, enabling effective multi-agent coordination.
@@ 3.5 Realistic Simulation Environments
Realistic simulation environments, such as OrbitZoo, must integrate high-fidelity physics models, flexible state representations, and scalable multi-agent frameworks. They should also provide interactive visualization tools and be publicly accessible to enable reproducibility and extensibility. By addressing the limitations of existing environments (e.g., lack of multi-agent support or simplified dynamics), OrbitZoo aims to advance learning-based approaches in space operations.
4 OrbitZoo: A Framework for Multi-Agent RL in Orbital Dynamics
Figure 1: High-level representation of OrbitZoo’s architecture. A JSON with information regarding a specific system must be provided to the environment (OrbitZoo). The environment then acts as the interface for generating data, developing single and multi-agent RL missions, or simply analyzing orbital dynamics.
OrbitZoo is a flexible and modular environment designed to address the challenges of applying multi-agent reinforcement learning (MARL) to high-fidelity orbital dynamics. It overcomes key limitations in existing tools, such as a lack of configurability, restricted support for realistic perturbative forces, and limited multi-agent capabilities. By integrating with Orekit, an industry-standard library for orbital mechanics, OrbitZoo ensures accuracy in physical modeling while remaining extensible for a wide range of RL missions. Its architecture, illustrated in Figure 1, provides modular components for data generation, mission design, and visualization, enabling researchers to systematically address the reality gap.
4.1 Architecture and Design
OrbitZoo’s architecture is designed to be modular and extensible. Figure 1 shows the primary modules, where gray boxes represent developed classes and other colors indicate integrated external components. This modular design allows users to implement and customize each component independently, ensuring compatibility with diverse RL algorithms and experimental setups.
4.2 Core Modules
Body: Modular Propagation of Orbital Dynamics
The Body class is the foundation of OrbitZoo and represents physical entities in the environment. Each body instance interfaces directly with Orekit’s numerical propagator, enabling high-fidelity simulations of orbital dynamics. The Body class incorporates: a unique identifier (alphanumeric value), a dry mass (in kg), a radius (in m), an expected Cartesian position and velocity $\mu=(x,y,z,\dot{x},\dot{y},\dot{z})$, and uncertainties associated with each of these elements $(\sigma_{x},\sigma_{y},\sigma_{z},\sigma_{\dot{x}},\sigma_{\dot{y}},\sigma_{\dot{z}})$, which are internally used to construct the covariance matrix $\Sigma$. Optionally, the initial date can be provided, enabling precise simulation of forces acting on the satellite at a specific moment. When resetting the environment, the actual position and velocity of the body are sampled from a multivariate normal distribution $\mathcal{N}(\mu,\Sigma)$. There are also static methods that provide information related to two given bodies, such as the distance between them, time of closest approach or probability of collision.
Satellite: Extending Bodies with Thrust Capabilities
The Satellite class extends the Body class to incorporate propulsion systems and agent parameters, essential for RL tasks requiring control and maneuverability. Satellites are configured with: (1) thrust parameters, including maximum thrust $(T_{\text{max}})$, deviation angle $(\theta_{\text{max}})$, and azimuthal angle $(\phi_{\text{max}})$ for realistic thrust modeling; (2) initial fuel mass and thrust-specific impulse for long-duration missions; (3) polar thrust parametrization, allowing satellites to apply thrust in any direction within physical constraints. This design enables RL agents to explore and optimize complex multi-dimensional action spaces; and (4) agent parameters. The agent parameters are arbitrary and serve as a standardization, depending on the specific requirements of the implemented algorithm for initialization.
Interface: Interactive Visualization
The Interface class provides interactive 3D visualization of the orbital environment, making OrbitZoo particularly useful for debugging, analysis, and presentation. It supports: (1) customizable visual components, such as equatorial grids, velocity and thrust vectors, and satellite trails; (2) real-time updates of system states, including timestamps, body names, and orbital parameters; and (3) flexible camera perspectives for inspecting multi-agent interactions and orbital trajectories.
(a) A system without bodies but with 3 different orbits.
(b) A system propagating several Starlink satellites.
(c) A single-agent mission of a satellite following a leader.
(d) A multi-agent mission of a constellation in GEO.
Figure 2: Frames of different systems on OrbitZoo’s interface.
Figure 2 demonstrates the visualization capabilities, ranging from single-agent missions to multi-agent constellations.
Additionally, we demonstrate this tool through four videos that highlight its capabilities in various scenarios. The first video offers an overview of the user interface, followed by demonstrations of the Hohmann maneuver mission, the GEO constellation mission, and a chase-target mission. These videos can be accessed through the following links: Interface Video; Hohmann Maneuver Mission; GEO Constellation Mission; Chase Target Mission.
The interface is built upon play3d, a library that leverages 2D perspective projections to create interactive 3D environments. Play3d is supported by the well-known 2D library pygame, which serves as its foundation.
Environment: High-Level Interaction
4.3 Addressing Key Challenges
OrbitZoo directly addresses the challenges outlined in Section 3:
Bridging the Reality Gap. OrbitZoo leverages Orekit’s high-fidelity propagator to model perturbative forces like atmospheric drag, solar radiation pressure, and third-body effects. A key feature is its flexibility in replicating real-world data, such as Starlink satellite orbits. This is enabled by its modular architecture, allowing precise customization of orbital parameters, perturbations, and thrust models. For instance, by optimizing parameters such as drag and reflection coefficients using Bayesian techniques, OrbitZoo can replicate Starlink’s orbital trajectories as provided by ephemeris data. This feature bridges the reality gap by ensuring that the simulated environment closely resembles the dynamics of real-world systems. As demonstrated in Section 5, the propagated trajectories generated by OrbitZoo achieve a low Root Mean Square Error (RMSE) when compared to Starlink ephemeris data, validating its suitability for high-fidelity reinforcement learning applications.
Realistic Thrust Modeling. The inclusion of polar thrust parametrization and configurable action spaces allows agents to perform realistic and efficient maneuvers, overcoming limitations in traditional thrust models.
Multi-Agent Coordination. OrbitZoo supports decentralized and centralized multi-agent training, enabling the simulation of satellite constellations and cooperative missions. Generalized advantage estimation (GAE) and PPO algorithms ensure stability and scalability in multi-agent setups.
Interactive Visualization for Debugging and Analysis. The visualization capabilities of OrbitZoo allow users to inspect agent behavior and system dynamics in real-time, making it essential for debugging and validating agent policies in a complex, partially observable environment.
@@ 5 Experiments and Results
To evaluate the effectiveness of OrbitZoo in modeling diverse orbital missions and supporting reinforcement learning (RL) methods, we conducted a series of experiments. These include a single-agent Hohmann transfer maneuver, a multi-agent geostationary orbit (GEO) constellation coordination problem, and a validation experiment comparing OrbitZoo’s simulations against real-world Starlink ephemeris data. These scenarios were chosen to demonstrate OrbitZoo’s ability to bridge the reality gap, enable realistic control tasks, and support scalable multi-agent coordination.
The main goal of these experiments was to validate OrbitZoo’s ability to model high-fidelity orbital dynamics, demonstrate the generalization of learned policies, and provide a robust environment for RL approaches, rather than aiming for state-of-the-art performance in each task. The following sections outline the setup and results of each experiment.
@@ 5.1 Single-Agent Hohmann Maneuver
The Hohmann transfer experiment was chosen as a benchmark to evaluate OrbitZoo’s ability to support reinforcement learning (RL) in a high-fidelity orbital dynamics environment. As a classic problem in orbital mechanics, the Hohmann transfer is analytically solvable and provides a clear reference for comparing RL-derived solutions with theoretical optima [22].
Setup
@@
Results
The RL agent learned near-optimal strategies for the Hohmann transfer, closely matching theoretical values for the semi-major axis. However, deviations in other orbital elements, such as inclination, occurred due to thrust misalignments. When tested with novel perturbative forces, the agent adapted and reached the target. These results validate OrbitZoo’s ability to model complex orbital dynamics and highlight opportunities to refine RL policy design and reward functions. Figure 12 shows the optimized transfer trajectory, achieving the target orbit with minimal fuel use. The trajectory closely matches the theoretical solution (Figure 3), demonstrating OrbitZoo’s high-fidelity modeling. Details of the experiment setup, training hyperparameters, and further quantitative results can be found in Appendix 0.D.2.
Figure 3: The L2 distance error between the optimal maneuver and Experiment 2 on the target orbit remains small over a large orbital period, with the larger error resulting from small inclination changes. However, the agents were trained to minimize differences in equinoctial elements rather than Euclidean distance.
@@ 5.2 Validation Against Real-World Data
To evaluate OrbitZoo’s ability to bridge the reality gap, we configured the environment to simulate Starlink satellites and compared its output to publicly available ephemeris data in Space-Track.
Setup
@@ Results
5.3 Multi-Agent GEO Constellation Coordination
This experiment demonstrates OrbitZoo’s ability to simulate and train multi-agent systems for satellite constellation management in geostationary orbit (GEO).
Setup
A four-satellite constellation in GEO aimed to maintain equal angular separation and altitude while minimizing fuel use through small thrusts. Each agent observes the anomalies of other satellites, its own semi-major axis and eccentricity, and outputs thrust commands via a decentralized policy. Training used PPO with generalized advantage estimation (GAE).
Results
Figure 5: Visualization of the initial random constellation and its evolution over nearly 4 days, with the purple circle representing the GEO orbit. After this period, the agents exhibit more distinct anomalies while maintaining proximity to the nominal orbit.
Figure 5 shows a view of the OrbitZoo interface of the constellation’s configuration after 4 days. The agents successfully maintained angular separation while minimizing fuel consumption. The policies are also able generalize to unseen perturbations, such as third body forces (Sun and Moon), solar radiation pressure and drag. Further results and a detailed explanation can be found in Appendix 0.D.5.
@@ 5.4 Additional Experiments
For completeness, we conducted a chase target experiment, where a satellite was tasked with pursuing a moving target in a higher orbit, and a single-agent collision avoidance experiment, where a satellite had to evade debris in LEO. The setup and results of these experiments are detailed in Appendix 0.D.3 and Appendix 0.D.4, respectively.
@@ Conclusions
We introduced OrbitZoo, a high-fidelity reinforcement learning (RL) environment for orbital dynamics that integrates with an industry-standard physics engine (Orekit). Unlike prior RL environments with simplified models, OrbitZoo incorporates realistic perturbative forces, flexible coordinate representations, and multi-agent support. It enables data-driven decision-making for space missions, bridging the reality gap between simulation and real-world orbital operations. Through a series of experiments—including a Hohmann transfer, a GEO constellation coordination task, and validation against real-world Starlink data—we demonstrated that RL policies trained in OrbitZoo can achieve near-optimal control strategies while aligning with physical satellite behavior. These results establish OrbitZoo as a credible benchmark for RL-based autonomy in space operations. Beyond RL, OrbitZoo serves as an open-source platform for researchers in aerospace engineering, space operations, and machine learning. Future work includes expanding real-world validation with more satellite datasets, integrating real-time onboard learning, and enhancing adaptive multi-agent coordination. By enabling realistic and reproducible RL research for space applications, OrbitZoo lays the groundwork for trustworthy AI-driven autonomy in future space missions.
Impact Statement
Reinforcement learning (RL) has shown promise in complex decision-making but struggles with real-world orbital dynamics due to the reality gap. OrbitZoo bridges this gap by providing a high-fidelity, modular RL simulation framework integrated with Orekit, supporting both single-agent and multi-agent tasks.
This work advances RL research in high-dimensional continuous control for space operations by incorporating perturbative forces, uncertainties, and ephemeris-based validation. Our experiments show that RL policies trained in OrbitZoo closely match real-world Starlink trajectories, establishing a benchmark for autonomous space systems.
Appendix 0.A Orbital Mechanics

### v2

@@ 3.3 Multi-Agent Coordination and Scalability
In MARL for orbital dynamics, agents operate in partially observable environments with decentralized policies. While MARL algorithms commonly assume fully cooperative scenarios (Dec-POMDP), in reality agents often assume partial or fully competitive scenarios (MA-POMDP). In such cases, task coordination, interference avoidance, and scaling to large constellations, that is, groups of satellites working together to accomplish shared objectives, become key challenges. EPyMARL papoudakis2021 extends PyMARL for flexible agent collaboration, while MARLlib MARLlib , compatible with PettingZoo environments like OrbitZoo, supports a broad set of algorithms and architectures. Federated Learning (FL) and Federated RL (FRL) FRL further enable decentralized learning and coordination in both cooperative and competitive settings while maintaining information privacy. Nonetheless, challenges inherent to the space environment – such as communication constraints, uncertainty, and heterogeneity between agents – make it challenging to apply standard MARL algorithms, highlighting the need for new, tailored approaches.
Recent solutions leverage C++ engines Zheng_2018 or CPU/GPU parallelization for faster training parallel-train and propagation parallel-step (see Appendix B.3). Scalability remains challenging when using high-fidelity numerical propagation, as each body experiences distinct forces that evolve continuously over time.
@@ 3.4 Realistic Simulation Environments
Realistic simulation environments should integrate (1) high-fidelity dynamics that account for the most relevant forces and perturbations (e.g., gravity fields, drag, SRP, third-body effects), (2) body and thruster specific characteristics, including shape, reflection and drag coefficients, and specific impulse, (3) flexible state and action spaces, as different missions may require distinct representations, (4) scalability and parallelization to handle an increasing number of bodies, and (5) reproducibility and extensibility. However, most existing MARL environments for orbital dynamics do not fully integrate these capabilities (Tab. 1).
4 Novel Reinforcement Learning Challenges in OrbitZoo
In the emerging era of autonomous systems operating in the physical world, RL is evolving beyond synthetic benchmarks and addressing real-world constraints such as actuation limits, uncertainty, safety, and coordination. OrbitZoo is designed to support this shift. It provides a modular testbed to rigorously investigate RL challenges grounded in high-fidelity dynamics while abstracting just enough to enable algorithmic insight. Below, we outline key research questions that OrbitZoo helps explore.
Sim-to-Real Transfer and Grounded Dynamics.
Sim-to-real gaps constitute a significant barrier in deploying RL policies in physical systems. While Orekit handles high-fidelity propagation, OrbitZoo provides the interfaces needed to test RL algorithms in scenarios with uncertain initial conditions, realistic gravity fields (Holmes–Featherstone harmonics holmes-featherstone ), drag (computed from historical weather data), SRP (accounting for the occlusion of the Sun by the Moon), and third-body forces (from all planets in the solar system, as well as the Sun, Moon, and the barycenters of the Earth–Moon system and the solar system). Validation against Starlink ephemerides (Sec. 5) shows that these perturbations can be tuned to match real trajectories, allowing research on robust policy learning under realistic dynamics.
Continuous Control and Thrust Constraints.
OrbitZoo supports physically plausible, time-continuous control through polar thrust parameterizations and constant-magnitude thrusts. Unlike many benchmarks that assume instantaneous or impulsive maneuvers, OrbitZoo models thrust as a continuous acceleration integrated over time, constrained by available fuel and propulsion limits. This formulation forces agents to learn realistic, sustained maneuvers, making the environment well suited for studying constrained RL and planning under resource limitations.
Multi-Agent Coordination with Partial Observability.
The environment supports decentralized, partially observable multi-agent setups via PettingZoo. These include tasks such as cooperative maneuvers, target chasing, and collision avoidance, scenarios rarely addressed in standard RL environments. OrbitZoo also enables centralized training and parallel simulations for scalability.
Reward Shaping under Sparse and Delayed Feedback.
Designing informative yet physically plausible reward signals is a key challenge in orbital control. OrbitZoo enables the use of parameterized reward functions, including orbital element targets (e.g., semi-major axis, inclination), fuel penalties, and safety-based margins, that enable flexible experimentation across shaped and sparse regimes. This supports research on credit assignment, reward misspecification, and curriculum learning in complex multi-agent environments.
Adversarial and Safety-Critical Learning.
OrbitZoo’s modular setup supports adversarial scenarios (e.g., pursuit-evasion, intentional jamming), allowing RL agents to be evaluated in high-stakes conditions. Combined with partial observability and fuel-aware dynamics, this supports research in safe, robust policy design.
Visualization and Debugging.
OrbitZoo offers a real-time, integrated 3D visualization tool that aids policy inspection and failure case diagnosis. Although not a core contribution, this feature supports reproducibility and interpretability, and is summarized in Appendix A.
@@ 5 Experiments and Results
To evaluate the effectiveness of OrbitZoo in modeling diverse orbital missions and supporting RL methods, we conducted a series of experiments using several learning algorithms. These include the development of a mission analogous to one created in a different environment, orbital transfers (OTs), collision avoidance maneuvers (CAMs), a multi-agent geostationary orbit (GEO) constellation coordination problem using independent and federated learning, and a validation experiment comparing OrbitZoo’s simulations against real-world Starlink ephemeris data. In the following missions, we consider an RL agent as a satellite with limited maneuvering capabilities, learning a strategy (policy) to apply thrusts that maximize the expected return (Sec. 3.1).
@@ 5.1 Single-Agent Hohmann Maneuver
The Hohmann transfer experiment was chosen as a benchmark to evaluate OrbitZoo’s ability to support RL in a high-fidelity orbital dynamics environment. Details of the experiment can be found in Appendix E.3. As a classic problem in orbital mechanics, the Hohmann transfer is analytically solvable and provides a clear reference for comparing RL-derived solutions with theoretical optima hohmann1960attainability .
Setup.
@@
Results.
The RL agent learned near-optimal strategies for a 30 km altitude Hohmann transfer, matching theoretical semi-major axis values. Deviations in other elements (e.g., inclination) arose from thrust misalignments. Contrary to the classical optimal maneuver, the agent adapted to new perturbations and reached the target successfully. These results validate OrbitZoo’s ability to model complex orbital dynamics and highlight areas for refining policy and reward design. Figure 15 shows the optimized transfer with minimal fuel use. The trajectory closely matches the theoretical solution (Figure 2), confirming OrbitZoo’s high-fidelity simulation. Full experimental details are presented in Appendix E.3.
5.2 Single-Agent Collision Avoidance Maneuver
CAM missions in RL are often focused on station-keeping while preventing collisions with other objects. Many existing approaches assume perfect knowledge of the current states of all bodies in the environment, simplifying the task to maximizing the Euclidean distance from other bodies while maintaining proximity to a nominal orbit. In reality, however, the states of these bodies are known with uncertainty, which must be accounted for in operational planning. In this mission, a more realistic approach to CAM is adopted by explicitly modeling state uncertainty and its evolution over time. While the environment handles propagation using high-fidelity dynamics that closely resemble real-world physics, the prediction of future states -— and the corresponding Probability of Collision (PoC) poc_akella -— relies on a simplified Newtonian attraction model, representing a low-fidelity simulation.
Setup.
Two bodies – debris and a satellite with maneuvering capabilities – are synthetically instantiated in the same Cartesian position with some uncertainty, and with symmetric Cartesian velocity, creating a short-term encounter at that moment. By propagating both bodies backward in time, we create the initial conditions for the start of each episode, enough for the satellite to perform the needed maneuvers and ultimately avoiding a collision. Due to the uncertainty, a PoC $10^{-}6$ is considered as high collision risk, and the agent is tasked at lowering it while staying near its initial orbit.
Results.
In many existing CAM approaches, the RL agent is trained using discrete action spaces, since maneuvers are only required when a collision risk exists. Accordingly, a DQN was employed for training the agent. To investigate whether applying thrust in specific directions could lead to improved performance, the continuous version of PPO was also implemented. In this setup, the action space included a decision variable ($0\leq\delta\leq 1$) indicating whether thrust should be applied ($\delta>0.5$). For training, the environment used Newtonian attraction with atmospheric drag to model the actual dynamics. During evaluation, however, all available perturbations in OrbitZoo were enabled to assess the generalization capability of the trained agents. Results show that both algorithms learned effective policies under the (simplified) training dynamics, but PPO demonstrated superior performance under realistic conditions, exhibiting a greater ability to reduce the PoC.
5.3 Multi-Agent GEO Constellation Coordination
This experiment demonstrates OrbitZoo’s ability to simulate and train multi-agent systems for satellite constellation management in geostationary orbit (GEO), using both independent and federated learning. Details can be found in Appendix E.6.
Setup.
A four-satellite constellation in GEO aimed to maintain equal angular separation and altitude while minimizing fuel use through small thrusts. Each agent observes the anomalies of other satellites, its own semi-major axis and eccentricity, and outputs thrust commands via a decentralized policy. Training used PPO with generalized advantage estimation (GAE) gae .
Results.
Figure 2 shows a view of the OrbitZoo interface of the constellation’s configuration after 4 days. The agents successfully maintained angular separation while minimizing fuel consumption. The policies are also able generalize to unseen perturbations, such as third body forces (Sun and Moon), SRP and drag. Further results and a detailed explanation can be found in Appendix E.6.
Figure 1: The L2 error between the optimal and Experiment 2 maneuvers stays low over long orbits, with most error due to minor inclination shifts. However, agents were trained to minimize equinoctial element differences, not Euclidean distance.
Figure 2: Visualization of the initial random constellation and its evolution over nearly 4 days, with the red circle representing the GEO orbit. After this period, the agents exhibit more distinct anomalies while maintaining proximity to the nominal orbit.
@@ 5.4 Validation Against Real-World Data
To evaluate OrbitZoo’s ability to bridge the reality gap, we configured the environment to simulate Starlink satellites and compared its output to publicly available ephemeris data in Space-Track. Further details on the analysis can be found in appendix G.
Setup.
@@ 5.5 Additional Experiments
For completeness, we present three additional experiments: (i) a mission implementation comparison with another environment (Appendix E.2), (ii) a chase-target scenario (Appendix E.4), where a satellite pursues a moving object in a higher orbit, and (iii) a single-agent collision avoidance task (Appendix E.5), where a satellite evades space debris in LEO. For each mission, we provide a brief overview, followed by the experimental setup, results, key challenges, and directions for future improvement.
@@ Conclusions
Through comparisons with existing environments and a diverse set of experiments—including a Hohmann transfer maneuver, a geostationary constellation coordination task, and validation using real-world Starlink ephemeris data—we demonstrated that RL policies trained in OrbitZoo can achieve near-optimal control while remaining consistent with realistic satellite behavior. These results establish OrbitZoo as a reliable benchmark for RL-based autonomy in space operations. To assess generalization and robustness, we employed multiple RL algorithms and different levels of realism, highlighting how continuous action spaces generally enhance performance under realistic orbital dynamics. Beyond RL, OrbitZoo serves as an open-source platform designed to support researchers in aerospace engineering, satellite operations, and machine learning. It provides a modular, high-fidelity environment for studying RL in realistic orbital settings. While grounded in orbital dynamics, its abstractions and challenge structures are representative of broader autonomy domains. The platform fosters reproducible experimentation and enables the exploration of autonomous decision-making for future space applications.
Appendix A OrbitZoo: A Framework for Multi-Agent RL in Orbital Dynamics
Figure 4: High-level overview of OrbitZoo’s architecture. A JSON file describing the orbital system is provided to OrbitZoo, which then serves as an interface for data generation, single- and multi-agent RL mission development, or orbital dynamics analysis.
OrbitZoo is a flexible and modular Python environment for RL designed to address the challenges of applying MARL to high-fidelity orbital dynamics. It overcomes key limitations in existing tools, such as a lack of configurability and visualization, restricted support for realistic perturbative forces, and limited multi-agent capabilities. While the vanilla OrbitZoo environment does not implement specific missions, RL algorithms, and training logic, it can easily be extended and configured for such cases.
A.1 Architecture and Design
OrbitZoo’s architecture is designed to be modular and extensible. Figure 4 shows the primary modules, where darker boxes represent developed classes and lighter colors indicate integrated external components. This modular design allows users to implement and customize each component independently, ensuring compatibility with diverse RL algorithms and experimental setups.
A.2 Core Modules
Body: Modular Propagation of Orbital Dynamics.
The Body class is the foundation of OrbitZoo and represents physical entities in the environment. Each body instance contains an individual numerical propagator, which is used to compute future states with great accuracy (position, velocity, uncertainty and mass) knowing the body’s current state and active forces. Upon initialization, each body must contain a unique identifier (name), a dry mass, a radius, an expected Cartesian position and velocity $\mu=(x,y,z,\dot{x},\dot{y},\dot{z})$, and uncertainties associated with each of these elements $(\sigma_{x},\sigma_{y},\sigma_{z},\sigma_{\dot{x}},\sigma_{\dot{y}},\sigma_{\dot{z}})$, which are internally used to construct the covariance matrix $\Sigma$ with these uncertainties as its diagonal entries. Optionally, the initial date can be provided, allowing an accurate representation of perturbative forces acting on the satellite at a specific moment. When resetting the environment, the actual position and velocity of the body are sampled from a multivariate normal distribution $\mathcal{N}(\mu,\Sigma)$. Some static methods also provide information related to two given bodies, such as the distance between them, time of closest approach, or probability of collision, which are useful for RL missions.
Satellite: Extending Bodies with Thrust Capabilities.
The Satellite class extends the Body class to incorporate propulsion systems and agent parameters, essential for RL tasks requiring control and maneuverability. Satellites are configured with: (1) polar thrust parametrization $(T,\theta,\phi)$ and action space, including maximum thrust $(T_{\text{max}})$, deviation angle $(\theta_{\text{max}})$, and azimuthal angle $(\phi_{\text{max}})$ for realistic thrust modeling; (2) initial fuel mass and thrust-specific impulse ($I_{\text{sp}}$) for long-duration missions; and (3) agent parameters. The agent parameters are arbitrary and serve as a standardization, depending on the specific requirements of the implemented algorithm for initialization.
Interface: Interactive Visualization.
The Interface class provides interactive 3D visualization of the orbital environment, making OrbitZoo particularly useful for debugging, analysis, and presentation. It supports: (1) customizable visual components, such as central body, equatorial grids, Keplerian orbits, velocity and thrust vectors, and satellite trails; (2) real-time updates of system states, including timestamps, body names, and orbital parameters; and (3) flexible camera perspectives for inspecting orbital trajectories and multi-agent interactions. Figure 5 demonstrates some visualization capabilities, ranging from single-agent missions to multi-agent constellations.
(a) A system with 3 different Keplerian orbits.
(b) A system propagating several Starlink satellites.
(c) A single-agent mission: Follow Leader.
(d) A multi-agent mission: Constellation in GEO.
Figure 5: Frames of different systems on OrbitZoo’s interface.
Additionally, we demonstrate this tool through four videos that highlight its capabilities in various scenarios. The first video offers an overview of the user interface, followed by demonstrations of the Hohmann maneuver mission, the GEO constellation mission, and a chase-target mission. These videos can be accessed through the following links: Interface Video; Hohmann Maneuver Mission; GEO Constellation Mission; Chase Target Mission. The interface is built upon play3d [1], a library that leverages 2D perspective projections to create interactive 3D environments.
Environment: High-level Interaction.
Appendix B Orbital Mechanics